OpenStudy (anonymous):

The Danville Express travels 280 mi at a certain speed. If the speed had been increased by 5 mph, the trip could have been made in 1 hr less time. Find the actual speed.

OpenStudy (nubeer):

speed =x, distance =280 time =t speed =x+5. distance= 280, time =t-1 distance = speed * time. try solving this way but not sure.

OpenStudy (anonymous):

Use the method of nubeer he is correct